Suppose the interest rate is 9% APR with monthly compounding. What is the present value of an annuity that pays $250 every three months for the next five years closest to? Convert APR to EAR and then use Texas Instruments BA II to solve using:

N:

I/Y:

PV:

PMT:

FV:

The present value of an annuity that pays $250 every three months for the next five years is closest to

A) $2280

B) $3985

C) $3990

D) $3995
Answer #1

1). Monthly rate = 9%/12 = 0.75%

EAR = [(1+0.75%)^12] -1 = 9.38%

This EAR compounded every 3 months will have a 3-monthly rate of [(1+9.38%)^(1/4)] -1 = 2.267%

N = 4*5 = 20; I/Y = 2.267%, FV = 0; PMT = -250, solve for PV.

PV = 3,984.49 (Option B)